Find the number of Na+ ions in 2.92 g of Na3PO4


1
Expert's answer
2022-02-14T17:01:06-0500

Solution:

The molar mass of Na3PO4 is 163.94 g/mol

Hence,

Moles of Na3PO4 = (2.92 g Na3PO4) × (1 mol Na3PO4 / 163.94 g Na3PO4) = 0.0178 mol Na3PO4


Na3PO4 → 3Na+ + PO43-

Hence,

1 mol of Na3PO4 contains 3 mol of Na+ ions

Thus, 0.0178 mol of Na3PO4 contains:

(0.0178 mol Na3PO4) × (3 mol Na+ ions / 1 mol Na3PO4) = 0.0534 mol Na+ ions


One mole of any substance contains 6.022×1023 particles.

Hence,

Number of Na+ ions = (0.0534 mol) × (6.022×1023 ions / 1 mol) = 3.216×1022 ions


Answer: 3.216×1022 of Na+ ions